OpenStudy (anonymous):

what is the equation of the line with slope -2 passing through (3,4)

OpenStudy (jadeishere):

y = -2x + 4

OpenStudy (jadeishere):

if you put it into slope-intercept form y = mx + b where y = b and m = the slope

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1\\ &({\color{red}{ 3}}\quad ,&{\color{blue}{ 4}})\quad \end{array} \\\quad \\ slope = {\color{green}{ m}}= -2 \\ \quad \\ y-{\color{blue}{ y_1}}={\color{green}{ m}}(x-{\color{red}{ x_1}})\qquad \textit{plug in the values and solve for "y"}\\ \qquad \uparrow\\ \textit{point-slope form} \)

OpenStudy (anonymous):

the only options to this answer is: y=-2x-10 y=-1/2x -5 y=-2x+10 7=2x-10

OpenStudy (jadeishere):

Then I don't know, i'm not in the right mind-set right now, sorry